# Every semi-normalized unconditional Schauder frame in Hilbert spaces contains a frame

### 11 Mar 2026 v2

@@ 1. Introduction
Let $g$ be a nonzero function in the Feichtinger algebra $M^{1}(\mathbb{R}^{d})$. Suppose that $M\subseteq L^{2}(\mathbb{R}^{d})$ is a closed subspace containing $\{e^{2\pi ib\cdot x}g(x)\}_{n\in\Gamma}$ for some nonzero $b\in\mathbb{R}^{d}$ and some uniformly discrete subset $\Gamma\subseteq\mathbb{R}^{d}$. Then $M$ does not admit any unconditional Schauder frame consisting of translates of finitely many functions.  $\diamondsuit$
@@ 2. Preliminaries
We say a linear operator $T\colon H\rightarrow H$ is positive (semi-definite) if $\langle\,Tx,x\,\rangle\geq 0$ for all $x\in H.$ A positive trace-class operator $T$ on $H$ is a positive operator on $H$ for which
@@ 3. Main Results
$\sum_{j=1}^{L_{n}}2^{-\ell_{n,j}}+\sum_{j=1}^{M_{n}}2^{-m_{n,j}}=\Big\lceil\sum_{j=1}^{L_{n}}2^{-\ell_{n,j}}\Big\rceil\quad\text{for all }n\in I.$
Let $\beta\in\mathbb{N}$ be such that $1<2^{\beta}\frac{\epsilon^{2}}{4C^{2}\delta}\leq 2$, where $\lceil\cdot\rceil$ is the ceiling function and $C>0$ is the absolute constant described in Theorem 2.2. We then define $\eta=\max\limits_{n,j}(\ell_{n,j},m_{n,j},\beta)$ and a family of positive trace-class operators $\{\phi_{n}\}_{n\in\mathbb{N}}$ that satisfies
@@
    $\text{tr}(\phi_{n})\,\leq\,2^{-\beta+2}\max\{\epsilon,\gamma\}.$
For each $n\in I,1\leq j\leq L_{n}$ we decompose $2^{-\ell_{n,j}}T_{n}$ and $2^{-m_{n,j}}\phi_{n}$ into
@@
Let $N=\eta-\beta$. We construct binary selectors of order $N$ as follows. Note that $2^{-N}\bigl(\,\#I_{1}+\#I_{2}\,\bigr)\in\mathbb{N}$. Consequently, it is guaranteed that we execute such a binary selecting process at least $N$ time. At first, we partition $I_{1}$ into sets of size $2$ by pairing $(n,i)$ that belong to $I_{1}$. For those $(n,i)\in I_{1}$ that is left unmatched, we randomly pair it with $(n,j)\in I_{2}.$ Next, we randomly pair the rest of $(n,j)\in I_{2}.$ Let $B_{b}$ be an arbitrary binary selector of order $1$. We partition $B_{b}$ with the same logic and then extract the binary selectors of order $2$. We then repeat this procedure $N$ times and obtain binary selectors of order $N.$ Note that for each $k\in I$ we have
@@
We are now ready to prove our main theorem as follows. It is worth noting that unconditional Schauder frames are sometimes defined as a pair of sequences such that Eqaution (1.3) holds with the unconditional convergence of the series in the norm of $H$ for all $x\in H$ in literature. The advantage of formulating unconditional Schauder frames in terms of a single sequence is that it allows one to establish the following equivalence between unconditional Schauder frames and sequences that can rescaled to form a frame for $H$ via Theorem 2.4.
@@
$x\,=\,\sum_{k\in I}\langle\,x,y_{n_{k}}\,\rangle\frac{x_{n_{k}}}{\|\,x_{n_{k}}\,\|}\,=\,\sum_{n\in\mathbb{N}}\langle\,x,z_{n}\,\rangle\,x_{n}\quad\text{for all }x\in H,$
@@
Without loss of generality, we may assume that $x_{n_{k}}\neq 0$ for all $k\in\mathbb{N}$. By Theorem 2.4, there exists some sequence of sequence of scalars such that $\{c_{n,k}x_{n_{k}}\}_{n\in\Gamma_{n_{k}},\,k\in\mathbb{N}}.$ Now by Lemma 3.2, there exists a sampling function $\sigma\colon\mathbb{N}\rightarrow\mathbb{N}$ such that $\{x_{n_{\sigma(k)}}\}_{k\in\mathbb{N}}$ can be normalized to form a frame for $H.$ In particular,
@@
where $A,B$ are frame bounds of $\bigl\{c_{n,k}x_{k}\bigr\}_{k\in\mathbb{N}}$ and $C$ is the absolute constant described in Theorem 2.2. Then for each $k\in\mathbb{N}$ we have
@@
Let $M=\max(144C^{2}BA^{-2},64C^{4}B^{-2}).$ Statement (d) then follows from the inequality
$M^{-2}\sum_{k\in\mathbb{N}}\bigl|\,\bigl\langle\,x,\,\frac{x_{n_{\sigma(k)}}}{\|\,x_{n_{\sigma(k)}}\,\|}\,\bigr\rangle\,\bigr|^{2}\,\leq\,\sum_{k\in\text{Range}(\sigma)}\bigl|\,\bigl\langle\,x,\,\frac{x_{n_{k}}}{\|\,x_{n_{k}}\,\|}\,\bigr\rangle\,\bigr|^{2}\,\leq\,\sum_{k\in\mathbb{N}}\bigl|\,\bigl\langle\,x,\,\frac{x_{n_{\sigma(k)}}}{\|\,x_{n_{\sigma(k)}}\,\|}\,\bigr\rangle\,\bigr|^{2}.$
@@
(b) Theorem 3.3 (d) provides a clearer characterization of the scalars chosen to rescale unconditional Schauder frames into a frame than Theorem 2.4 when the notion of an unconditional Schauder frame is formulated as a single sequence. Specifically, for any unconditional Schauder frame $\{x_{n}\}_{n\in\mathbb{N}}$ for $H$ there always exists a subset $I\subseteq\mathbb{N}$ such that one can always rescale $\{x_{n}\}_{n\in\mathbb{N}}$ into a frame by using the rescaling scalars $c_{n}=\|\,x_{n}\,\|^{-1}$ if $n\in I$ and $c_{n}=0$ if $n\notin I$. Let $\{y_{n}\}_{n\in\mathbb{N}}$ be coefficient functionals associated with $\{x_{n}\}_{n\in\mathbb{N}}$, we remark that it remains unclear to us whether there exists some subset $J\subseteq\mathbb{N}$ such that both $\bigl\{\frac{x_{n}}{\|\,x_{n}\,\|}\bigr\}_{n\in J}$ and $\{\|\,x_{n}\,\|y_{n}\}_{n\in J}$ are frames for $H.$ By Orlicz’s Theorem ([47]), $\{\|\,x_{n}\,\|y_{n}\}_{n\in J}$ must be a Bessel sequence. However, the existence of a lower frame bound for $\bigl\{\|\,x_{n}\,\|y_{n}\bigr\}_{n\in J}$ remains unknown.  $\diamondsuit$
@@ 4.2. Beurling density associated with Gabor systems
Next, we construct an unconditional Schauder frame with infinite Beurling density for which all associated coefficient functionals are nonzero. Let $\phi,\Gamma$ be such that ${\mathcal{G}}(\phi,\Gamma)$ forms a frame for $L^{2}(\mathbb{R}^{d})$ and let $\{\phi^{*}_{n}\}_{n\in\mathbb{N}}$ be the canonical dual frame associated with ${\mathcal{G}}_{(}\phi,\Gamma).$ Note that we have $\sup_{n}\|\,\phi_{n}^{*}\,\|<\infty.$ Let $(K_{n})_{n\in\mathbb{N}}$ be an increasing sequence of positive integer such that $\sum_{n=1}^{\infty}\frac{1}{K_{n}}<\infty.$ For each $n\in\mathbb{N}$ we pick $K_{n}$ distinct points $\{(a_{n,i},b_{n,i})\}_{i=1}^{K_{n}}$ near $(a_{n},b_{n})$ such that
@@
We will show that $S$ is an invertible bounded linear operator. We will also show that the series in Equation 4.1 converges unconditionally in the norm of $L^{2}(\mathbb{R}^{d}).$ It then follows that
@@
with the unconditional convergence of the series in the norm of $L^{2}(\mathbb{R}^{d}).$ Since both sequences $\cup_{n\in\mathbb{N}}\{K^{-1/2}\phi_{n}^{*}\}_{i=1}^{K_{n}}$ and $\cup_{n\in\mathbb{N}}\{K^{-1/2}M_{b_{n}}T_{a_{n}}\phi\}_{i=1}^{K_{n}}$ are Bessel sequences, we see that
@@
with the unconditional convergence of the series in the norm of $L^{2}(\mathbb{R}^{d}).$ By the Triangle Inequality, for each $n\in\mathbb{N}$ we have
$\begin{split}\|\,\langle\,x,K_{n}^{-1}\phi_{n}^{*}\,\rangle M_{b_{n,i}}T_{a_{n,i}}\phi\,\|&\,\leq\,K_{n}^{-1}\bigl(\,\|\,\langle\,x,\phi_{n}^{*}\,\rangle(M_{b_{n,i}}T_{a_{n,i}}-M_{b_{n}}T_{a_{n}})\phi\,\|+\|\,\langle\,x,\phi_{n}^{*}\,\rangle M_{b_{n}}T_{a_{n}}\phi\,\|\,\bigr)\\ &\,\leq\,(K_{n}^{-1}2^{-2n}\sup_{n}\|\,\phi_{n}^{*}\,\|)\|\,x\,\|.\end{split}$
It follows that, for any positive integers $N<M$ and any $x\in H$ with $\|\,x\,\|=1$, we have
$\begin{split}\Bigl\|\,\sum_{n=N}^{M}\sum_{i=1}^{K_{n}}\langle\,x,K_{n}^{-1}\phi_{n}^{*}\,\rangle M_{b_{n,i}}T_{a_{n,i}}\phi\,\Bigr\|\,\leq\,\sup_{n}\|\,\phi_{n}^{*}\,\|\sum_{n=N}^{M}2^{-2n}+\Bigl\|\,\sum_{n=N}^{M}\sum_{i=1}^{K_{n}}\langle\,x,K_{n}^{-1}\phi_{n}^{*}\,\rangle M_{b_{n}}T_{a_{n}}\phi\,\Bigr\|.\end{split}$
Consequently, $S(x)$ is a well-defined linear operator. Using the Triangle Inequality again and arguing similarly, we see that $\|\,S-I\,\|<1$, and hence $S$ is invertible. Moreover, the series in Equation (4.1) converges unconditionally due to the unconditional convergence of the series in Equation (4.2). ∎
@@ 4.4. Exponential Systems
Landau’s necessary density conditions states that $D^{-}(\Lambda)\geq|S|$ whenever ${\mathcal{E}}(\Lambda)$ forms a frame for $L^{2}(S)$, and that $D^{+}(\Lambda)\leq|S|$ if ${\mathcal{E}}(\Lambda)$ forms a Riesz basis for its closed span. Here $|S|$ denotes the Lebesgue measure of $S.$ Consequently, $D^{-}(\Lambda)=D^{+}(\Lambda)=|S|$ if ${\mathcal{E}}(\Lambda)$ is a Riesz basis for $L^{2}(S).$ As in the case of Gabor systems, it remains unknown how small $D^{-}(\Lambda)$ can be when ${\mathcal{E}}(\Lambda)$ forms an unconditional Schauder frame for $L^{2}(S).$ By Theorem 3.3, we obtain the following necessary density condition for such exponential systems. We remark that, by an argument similar to that used in Theorem 4.7, one can also construct $\Lambda\subseteq\mathbb{R}^{d}$ with $D^{+}(\Lambda)=\infty$ such that ${\mathcal{E}}(\Lambda)$ forms an unconditional Schauder frame that possess coefficient functionals that are all nonzero.

### 13 Mar 2026 v3

@@ 1. Introduction
Let $g$ be a nonzero function in the Feichtinger algebra $M^{1}(\mathbb{R}^{d})$. Suppose that $M\subseteq L^{2}(\mathbb{R}^{d})$ is a closed subspace containing $\{e^{2\pi ib\cdot x}g(x)\}_{b\in\Gamma}$ for some infinite uniformly discrete subset $\Gamma\subseteq\mathbb{R}^{d}$. Then $M$ does not admit any unconditional Schauder frame consisting of translates of finitely many functions.  $\diamondsuit$
@@ 2. Preliminaries
We say that a linear operator $T\colon H\rightarrow H$ is positive (semi-definite) if $\langle\,Tx,x\,\rangle\geq 0$ for all $x\in H.$ A positive trace-class operator $T$ on $H$ is a positive operator on $H$ for which
@@ 3. Main Results
$\sum_{j=1}^{L_{n}}2^{-\ell_{n,j}}+\sum_{j=1}^{M_{n}}2^{-m_{n,j}}=\Big\lceil\sum_{j=1}^{L_{n}}2^{-\ell_{n,j}}\Big\rceil\quad\text{for all }n\in I,$
where $\lceil\cdot\rceil$ is the ceiling function. Let $\beta\in\mathbb{N}$ be such that $1<2^{\beta}\frac{\epsilon^{2}}{4C^{2}\delta}\leq 2$ where $C>0$ is the absolute constant described in Theorem 2.2. We then define $\eta=\max\limits_{n,j}(\ell_{n,j},m_{n,j},\beta)$ and a family of positive trace-class operators $\{\phi_{n}\}_{n\in I}$ that satisfies
@@
    $\text{tr}(\phi_{n})\,\leq\,2^{-\beta+2}\max\{\epsilon,\gamma\},$
for all $n\in I$. For each $n\in I,1\leq j\leq L_{n}$ we decompose $2^{-\ell_{n,j}}T_{n}$ and $2^{-m_{n,j}}\phi_{n}$ into
@@
Let $N=\eta-\beta$. We construct binary selectors of order $N$ as follows. Note that $2^{-N}\bigl(\,\#I_{1}+\#I_{2}\,\bigr)\in\mathbb{N}$. Consequently, it is guaranteed that we execute such a binary selecting process at least $N$ time. At first, we partition $I_{1}$ into sets of size $2$ by pairing $(n,i)$ that belong to $I_{1}$ and share the same first index $n$. For those $(n,i)\in I_{1}$ that is left unmatched, we randomly pair it with $(n,j)\in I_{2}.$ Next, we randomly pair the rest of $(n,j)\in I_{2}.$ Let $B_{b}$ be an arbitrary binary selector of order $1$. We partition $B_{b}$ with the same logic and then extract the binary selectors of order $2$. We then repeat this procedure $N$ times and obtain binary selectors of order $N.$ Note that for each $k\in I$ we have
@@
We are now ready to prove our main theorem as follows. It is worth noting that unconditional Schauder frames are sometimes defined as a pair of sequences such that Eqaution (1.3) holds with the unconditional convergence of the series in the norm of $H$ for all $x\in H$ in literature. The advantage of formulating unconditional Schauder frames in terms of a single sequence is that it allows one to establish the following equivalence between unconditional Schauder frames and sequences that can be rescaled to form a frame for $H$ via Theorem 2.4.
@@
$x\,=\,\sum_{k\in I}\langle\,x,y_{k}\,\rangle\frac{x_{n_{k}}}{\|\,x_{n_{k}}\,\|}\,=\,\sum_{k\in\mathbb{N}}\langle\,x,z_{k}\,\rangle\,x_{k}\quad\text{for all }x\in H,$
@@
Without loss of generality, we may assume that $x_{n_{k}}\neq 0$ for all $k\in\mathbb{N}$. By Theorem 2.4, there exists some sequence of scalars such that $\{c_{n,k}x_{n_{k}}\}_{n\in\Gamma_{n_{k}},\,k\in\mathbb{N}}.$ Now by Lemma 3.2, there exists a sampling function $\sigma\colon\mathbb{N}\rightarrow\mathbb{N}$ such that $\{x_{n_{\sigma(k)}}\}_{k\in\mathbb{N}}$ can be normalized to form a frame for $H.$ In particular,
@@
where $A,B$ are frame bounds of $\bigl\{c_{n,k}x_{n_{k}}\bigr\}_{k\in\mathbb{N}}$ and $C$ is the absolute constant described in Theorem 2.2. Then for each $k\in\mathbb{N}$ we have
@@
Let $L=\max(144C^{2}BA^{-2},64C^{4}B^{-2}).$ Statement (d) then follows from the inequality
$L^{-2}\sum_{k\in\mathbb{N}}\bigl|\,\bigl\langle\,x,\,\frac{x_{n_{\sigma(k)}}}{\|\,x_{n_{\sigma(k)}}\,\|}\,\bigr\rangle\,\bigr|^{2}\,\leq\,\sum_{k\in\text{Range}(\sigma)}\bigl|\,\bigl\langle\,x,\,\frac{x_{n_{k}}}{\|\,x_{n_{k}}\,\|}\,\bigr\rangle\,\bigr|^{2}\,\leq\,\sum_{k\in\mathbb{N}}\bigl|\,\bigl\langle\,x,\,\frac{x_{n_{\sigma(k)}}}{\|\,x_{n_{\sigma(k)}}\,\|}\,\bigr\rangle\,\bigr|^{2}.$
@@
(b) Theorem 3.3 (d) provides a clearer characterization of the scalars chosen to rescale unconditional Schauder frames into a frame than Theorem 2.4 when the notion of an unconditional Schauder frame is formulated as a single sequence. Specifically, for any unconditional Schauder frame $\{x_{n}\}_{n\in\mathbb{N}}$ for $H$ there always exists a subset $I\subseteq\mathbb{N}$ such that one can always rescale $\{x_{n}\}_{n\in\mathbb{N}}$ into a frame by using the rescaling scalars $c_{n}=\|\,x_{n}\,\|^{-1}$ if $n\in I$ and $c_{n}=0$ if $n\notin I$. Let $\{y_{n}\}_{n\in\mathbb{N}}$ be the sequence of coefficient functionals associated with $\{x_{n}\}_{n\in\mathbb{N}}$, we remark that it remains unclear to us whether there exists some subset $J\subseteq\mathbb{N}$ such that both $\bigl\{\frac{x_{n}}{\|\,x_{n}\,\|}\bigr\}_{n\in J}$ and $\{\|\,x_{n}\,\|y_{n}\}_{n\in J}$ are frames for $H.$ By Orlicz’s Theorem ([47]), $\{\|\,x_{n}\,\|y_{n}\}_{n\in J}$ must be a Bessel sequence. However, the existence of a lower frame bound for $\bigl\{\|\,x_{n}\,\|y_{n}\bigr\}_{n\in J}$ remains unknown.  $\diamondsuit$
@@ 4.1. Olson-Zalik Conjecture
Using the fact that modulation spaces are invariant under Fourier transform, we obtain the following corollary.
@@
We provide an illustrative example below how Theorem 4.2 improves the results in [17], [34] and [52].
@@ 4.2. Beurling density associated with Gabor systems
Next, we construct an unconditional Schauder frame with infinite Beurling density for which all associated coefficient functionals are nonzero. Let $\phi,\Gamma$ be such that ${\mathcal{G}}(\phi,\Gamma)$ forms a frame for $L^{2}(\mathbb{R}^{d})$ and let $\{\phi^{*}_{n}\}_{n\in\mathbb{N}}$ be the canonical dual frame associated with ${\mathcal{G}}_{(}\phi,\Gamma).$ Note that we have $\sup_{n}\|\,\phi_{n}^{*}\,\|_{L^{2}(\mathbb{R}^{d})}<\infty.$ Let $(K_{n})_{n\in\mathbb{N}}$ be an increasing sequence of positive integer such that $\sum_{n=1}^{\infty}\frac{1}{K_{n}}<\infty.$ For each $n\in\mathbb{N}$ we pick $K_{n}$ distinct points $\{(a_{n,i},b_{n,i})\}_{i=1}^{K_{n}}$ near $(a_{n},b_{n})$ such that
@@
We will show in the following that $S$ is an invertible bounded linear operator. We will also show that the series in Equation 4.1 converges unconditionally in the norm of $L^{2}(\mathbb{R}^{d}).$ Once these two claims are established, it then follows that
@@
with the unconditional convergence of the series in the norm of $L^{2}(\mathbb{R}^{d}).$ Consequently, $\bigl\{M_{b_{n,i}}T_{a_{n,i}}\phi\,|\,1\leq i\leq K_{n},\,n\in\mathbb{N}\bigr\}$ is an unconditional Schauder frame for $L^{2}(\mathbb{R}^{d}).$
Since both sequences $\cup_{n\in\mathbb{N}}\{K^{-1/2}\phi_{n}^{*}\}_{i=1}^{K_{n}}$ and $\cup_{n\in\mathbb{N}}\{K^{-1/2}M_{b_{n}}T_{a_{n}}\phi\}_{i=1}^{K_{n}}$ are Bessel sequences, we see that
@@
with the unconditional convergence of the series in the norm of $L^{2}(\mathbb{R}^{d}).$ By applying condition (c) above and the Cauchy–Bunyakovsky–Schwarz inequality, for each $n\in\mathbb{N}$ we obtain
$\begin{split}\|\,\langle\,x,\phi_{n}^{*}\,\rangle(M_{b_{n,i}}T_{a_{n,i}}-M_{b_{n}}T_{a_{n}})\phi\,\|_{L^{2}(\mathbb{R}^{d})}&\,\leq\,\sup_{n}\|\,\phi_{n}^{*}\,\|_{L^{2}(\mathbb{R}^{d})}\|\,x\,\|_{L^{2}(\mathbb{R}^{d})}(2^{2n}\sup_{n}\|\,\phi_{n}^{*}\,\|_{L^{2}(\mathbb{R}^{d})})^{-1}\\ &\,\leq\,2^{-2n}\|\,x\,\|_{L^{2}(\mathbb{R}^{d})}.\end{split}$ (4.3)
By the identity $M_{b_{n,i}}T_{a_{n,i}}\phi=(M_{b_{n,i}}T_{a_{n,i}}\phi-M_{b_{n}}T_{a_{n}}\phi)+M_{b_{n}}T_{a_{n}}\phi$, Equation (4.3) and Triangle Inequality, it follows that, for any positive integers $N<M$ and any $x\in L^{2}(\mathbb{R}^{d})$ with $\|\,x\,\|_{L^{2}(\mathbb{R}^{d})}=1$, we have
$\begin{split}\Bigl\|\,\sum_{n=N}^{M}\sum_{i=1}^{K_{n}}\langle\,x,K_{n}^{-1}\phi_{n}^{*}\,\rangle M_{b_{n,i}}T_{a_{n,i}}\phi\,\Bigr\|_{L^{2}(\mathbb{R}^{d})}\,\leq\,\sum_{n=N}^{M}2^{-2n}+\Bigl\|\,\sum_{n=N}^{M}\sum_{i=1}^{K_{n}}\langle\,x,K_{n}^{-1}\phi_{n}^{*}\,\rangle M_{b_{n}}T_{a_{n}}\phi\,\Bigr\|_{L^{2}(\mathbb{R}^{d})}.\end{split}$
Consequently, $S(x)$ is a well-defined linear operator. Using the Triangle Inequality again and arguing similarly, we see that
$\|\,(S-I)x\,\|_{L^{2}(\mathbb{R}^{d})}=\Bigl\|\,\sum_{n\in\mathbb{N}}\sum_{i=1}^{K_{n}}\langle\,x,K_{n}^{-1}\phi_{n}^{*}\,\rangle M_{b_{n,i}}T_{a_{n,i}}\phi-\sum_{n=1}^{\infty}\langle\,x,\phi_{n}^{*}\,\rangle M_{b_{n}}T_{a_{n}}\phi\,\Bigr\|_{L^{2}(\mathbb{R}^{d})}<1,$
for all $x\in L^{2}(\mathbb{R}^{d})$ with $\|\,x\,\|_{L^{2}(\mathbb{R}^{d})}=1$. Hence, $S$ is an invertible bounded linear operator. Moreover, the series in Equation (4.1) converges unconditionally due to the unconditional convergence of the series in Equation (4.2). ∎
@@ 4.4. Exponential Systems
Landau’s necessary density conditions states that $D^{-}(\Lambda)\geq|S|$ whenever ${\mathcal{E}}(\Lambda)$ forms a frame for $L^{2}(S)$, and that $D^{+}(\Lambda)\leq|S|$ if ${\mathcal{E}}(\Lambda)$ forms a Riesz basis for its closed span. Consequently, $D^{-}(\Lambda)=D^{+}(\Lambda)=|S|$ if ${\mathcal{E}}(\Lambda)$ is a Riesz basis for $L^{2}(S).$ As in the case of Gabor systems, it remains unknown how small $D^{-}(\Lambda)$ can be when ${\mathcal{E}}(\Lambda)$ forms an unconditional Schauder frame for $L^{2}(S).$ By Theorem 3.3, we obtain the following necessary density condition for such exponential systems. We remark that, by an argument similar to that used in Theorem 4.7, one can also construct $\Lambda\subseteq\mathbb{R}^{d}$ with $D^{+}(\Lambda)=\infty$ such that ${\mathcal{E}}(\Lambda)$ forms an unconditional Schauder frame that possess coefficient functionals that are all nonzero.